In a random sample of 19 residents of the state of Texas, the mean waste recycled per person per day was 2.5 pounds with a standard deviation of 0.650.65 pounds. Determine the 95% confidence interval for the mean waste recycled per person per day for the population of Texas. Assume the population is approximately normal.

Step 1 of 2:

Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places.
Answer #1

Solution :

degrees of freedom = n - 1 = 19 - 1 = 18

t/2,df = t0.025,18 = 2.101

Margin of error = E = t/2,df * (s /n)

= 2.101 * ( 0.65/ 19)

Margin of error = E = 0.313

The 95% confidence interval estimate of the population mean is,

- E < < + E

2.5 - 0.313 < < 2.5 + 0.313

( 2.187 < < 2.813 )